Write the equation in slope intercept form of the line perpendicular to the graph
Y= -3/4x + 7

Answer by mananth(16946)

1 y = - 3/ 4 x + 7

Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = - 3/4

The slope of a line perpendicular to the above line will be the negative reciprocal 4/3
Because m1*m2 =-1
The slope of the required line will be 4/3
So the equation will be y = 4/3 x +c